# Accuracy Assessment of OpenAlex and Clarivate Scholar ID with an LLM-Assisted Benchmark

### v1

@@ 2.3.1 Disambiguation Accuracy and Dataset Integrity
Post-disambiguation analysis yielded 523 Chinese authors, 241 American authors, and 269 European authors. We capped each regional activity level module at 100 authors, supplementing with uniquely email-matched authors when necessary.
The final standard datasets comprised 900 authors each for China and the USA, while Europe’s smaller pool included 766 authors. Consequently, our regional comparative analysis focused primarily on China and the USA.
@@ 3.2 Recall and Precision of OpenAlex ID
To evaluate the precision and recall of author identification in OpenAlex, we conducted experiments using a dataset of Chinese scholars. The standard dataset was selected from Clarivate Analytics, focusing on papers published in the top quartile (Q1) journals within the last five years. This standard set comprised 900 authors.
First, we obtained the home page address (author_id) for each author by traversing all their corresponding papers. If the author_id could not be retrieved through their papers, we searched for the author’s name combined with their institutional information to obtain the author_id. Using these author_ids, we collected all papers where the author was listed as the corresponding author and had published in Q1 journals.
In our dataset of 900 authors, 102 were found to have multiple author_ids, 100 did not have any corresponding author_id, and 182 authors could not be linked to any paper where they were the corresponding author.
We then intersected the standard dataset and the papers retrieved from OpenAlex, resulting in 463 common papers. We extracted detailed information for these papers from both Clarivate Analytics and OpenAlex, including publication year, number of authors, the position of the scholar in the author list, and institutional affiliation.
Upon comparison, we discovered discrepancies in the publication years for 162 papers between Clarivate Analytics and OpenAlex. Verification through Google Scholar confirmed that the publication years provided by Clarivate Analytics were accurate, indicating significant errors in the basic information of numerous papers in OpenAlex.
Additionally, during the experiment, we observed indexing errors in OpenAlex. Specifically, an author’s paper appeared in the list of papers on his/her profile page, but it was missing when retrieved through the API.
These findings highlight critical issues in the accuracy and completeness of data in OpenAlex, suggesting a need for further improvement in its data curation processes.
@@
Data and code, including prompts used in build the agents are available at following github project:

### v2

@@ 2.3.1 Disambiguation Accuracy and Dataset Integrity
Post-disambiguation analysis yielded 523 Chinese authors, 241 American authors, and 269 European authors. We capped each regional activity level module at 100 authors, supplementing with uniquely email-matched authors when necessary. The final standard datasets comprised 900 authors each for China, the USA and Europe.
@@ 3.1 Recall and Precision of Clarivate ID
Figure 6: Precision and Recall of OpenAlex ID in China.
Figure 7: Precision and Recall of OpenAlex ID in USA.
Figure 8: Precision and Recall of OpenAlex ID in Europe.
@@ 3.2 Recall and Precision of OpenAlex ID
In this section, we evaluate the OpenAlex dataset using a standard dataset of 900 Chinese scholars whose work has been published in the first quartile over the past five years according to Clarivate Analytics. The evaluation process involves several steps as detailed below.
Firstly, we traverse all papers corresponding to each author to obtain their homepage addresses, also known as author_id. If the author_id cannot be retrieved through the papers, we search for the scholar’s name along with their institutional information to acquire the corresponding author_id. Using the author_id, we then retrieve all their papers from OpenAlex, filtering only those published in the first quartile where the author is the corresponding author.
Due to inconsistent formatting of paper titles across databases, many exact matches might fail because of extraneous symbols. The average length of the paper titles was found to be 108 characters. In this experiment, we used an edit distance of 15; if the edit distance between two titles is less than or equal to 15, they are considered the same.
We employed two methods to find the intersection of papers. First, all papers obtainable through the OpenAlex API were included in the intersection, which resulted in 851 scholars having non-empty paper sets in OpenAlex. Second, the intersection was taken between the standard dataset and the papers retrieved via author_id, resulting in 657 scholars with non-empty intersections. It should be noted that some papers in OpenAlex can be queried but do not appear on the author’s homepage due to incorrect disambiguation or classification. Therefore, the first method is considered more representative.
Issues were identified with the OpenAlex dataset:
  1. 1.
    Taking the intersection of papers obtained from the standard dataset and OpenAlex, 463 papers were analyzed for detailed information such as publication year, number of authors, author’s position in the author list, and affiliated institutions. Discrepancies in publication year were found in 162 papers when compared to Clarivate Analytics, verified to be correct by Google Scholar and other publication databases, indicating significant errors in basic information within OpenAlex.
  2. 2.
    An unexpected indexing error in OpenAlex was discovered: a paper (let’s call it ’a’) listed on an author’s homepage could not be retrieved via the API.
  3. 3.
    We also found inadequacies in the disambiguation process. Identical papers appeared under multiple paper_ids. For instance, the paper titled "Glycosyltransferase UGT79B7 negatively regulates hypoxia response through $\gamma$aminobutyric acid homeostasis in Arabidopsis" had two results in OpenAlex: https://openalex.org/W3135641188 and https://openalex.org/W3194753228.
Among the 900 Chinese scholars, 851 had non-empty intersections. 87 scholars had multiple author_id links, 72 lacked author_id, and 176 experienced multiple returns for the same title. The most frequent case showed 8 out of 23 papers being duplicated. For 900 American scholars, 869 had non-empty intersections, 54 had multiple author_ids, 108 lacked author_ids, and 391 experienced multiple returns for the same title.
Figure 9: Precision of OpenAlex ID.
Figure 10: Recall of OpenAlex ID.
As shown in Figure 9 and Figure 10, the precision across various regions is not high, indicating challenges faced by OpenAlex in accurately disambiguating scholars with identical names. When extracting all institutions for a single scholar, if a scholar has affiliations a and b for paper A, and a and c for paper B, we consider a, b, and c as belonging to the same author. For example, under "zhang, zhiguo", there are three unrelated institution sets, likely indicating three different individuals with the same name.
Additionally, a combined comparison among three regions reveals that the recall rate in the information electronics field exceeds 90%, significantly higher than in the other two fields. This could be attributed to broader collaboration and more active research in the information electronics domain, leading to better classification outcomes.
@@
Data and code will be available at github.com/sylviachen0513.